pur31024:

A plastic bin contains red, yellow, and green key chains. Out of 350 key chains, 10% percent are yellow. Of the remaining key chains, are green. How many red key chains are inside the bin? A. 26 B. 35 C. 126 D. 189

russkaya05:

Red+green+yellow=350 10% out of 350 are yellow 0.1 times 350=35 350-35=315 3/5 out of 315 are green red is not green so 2/5 of what is left is red 2/5 of 315 is red 2/5 times 315=430/5=126 126 red try this please i think thats how it works

russkaya05:

text me if it is correct

Hero:

And how do you know how many are green @russkaya05

russkaya05:

I just explained it on top

russkaya05:

\[3/5\] out of 315 are green

Hero:

The given information only says 10% are yellow. It doesn't say how many are green. @russkaya05